OpenStudy (asylum15):

- Can someone help me out with a mechanics uncertainty? Just a clarification. PLEASE!

OpenStudy (asylum15):

\[I = \frac{ \pi(D^4) }{ 64 }\]

OpenStudy (asylum15):

What dictates the number under the line? 64 = a cylinder? 32 = ?

OpenStudy (anonymous):

D hope fully stand s for Diameter and radius=D/2 u have written I=(pi)/4 (D/2)^4 (may be i am not sure as exactly in what context u need it)

OpenStudy (asylum15):

Pi by D to the power of 4 OVER 64. This finds I, a moment of inertia. This is part of the formula, M/I = Stress/Y = E/R An engineering formula

OpenStudy (anonymous):

yup that much i guessed too.. thats why i hope D stands for Diameter

OpenStudy (asylum15):

So, any ideas?

OpenStudy (anonymous):

are u finding the MI of a circular disc

OpenStudy (asylum15):

Of a human femur

OpenStudy (anonymous):

than i may not be correct for D as diameter

OpenStudy (asylum15):

The femur is taken as cylinderical in this question.

OpenStudy (asylum15):

The D doesn't matter though does it?

OpenStudy (anonymous):

D matters and usually diameter in mechanics..

OpenStudy (anonymous):

may i know from which book u r referring to

OpenStudy (asylum15):

These are notes given to me by my lecturer. I believe that 64 is used when the shape being studied is a cylinder. If a beam or irregular shape, it becomes 32 I believe..